A helium balloon (in the shape of a sphere) has radius 7.10 m .

For the density of air, please use 1.29 kg/m^3, and for Helium, use .179 kg/m^3

Part A How much additional mass (payload) could th
Answer #1

Mass of air displaced by balloon = 4/3 * pi * 7.13 x 1.29kg/m^3 = 1934kg = Mass of loaded balloon

Mass of unloaded balloon = 930kg + 4/3 * pi * 7.13 x 0.179 kg/m^3 = 1198kg

Mass that balloon can lift = Mass of loaded balloon - Mass of unloaded balloon = 1934kg - 1198kg = 736kg